HIGH-RESOLUTION SPECTROSCOPY OF THE ν_{8} BAND OF METHYLENE BROMIDE USING A QUAN-TUM CASCADE LASER-BASED CAVITY RINGDOWN SPECTROMETER

\underline{JACOB T. STEWART}, BRIAN E. BRUMFIELD, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana, IL 61801; MATTHEW D. ESCARRA, CLAIRE F. GMACHL, Department of Electrical Engineering, Princeton University, Princeton Institute for the Science and Technology of Materials, Princeton, NJ 08544; BENJAMIN J. McCALL, Departments of Chemistry and Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL 61801.

Our group has constructed a cavity ringdown spectrometer based on a quantum cascade laser (QCL) in an effort to obtain a high-resolution gas phase spectrum of buckminsterfullerene (C_{60}). To test the performance of our spectrometer we have observed the ν_{8} band of methylene bromide (CH_{2}Br_{2}) from 1196 to 1197.5 cm^{−1}. This band had previously only been recorded at low resolution. Cold methylene bromide molecules were produced in a continuous supersonic expansion from a 700 μm pinhole and probed using continuous wave cavity ringdown spectroscopy (cw-CRDS). To our knowledge, this is the first experiment to measure a supersonic jet-cooled sample using a QCL-based cw-CRDS spectrometer. We have assigned the observed ro-vibrational lines from the three isotopomers of CH_{2}Br_{2} to effective Hamiltonians, and find that the molecules have been cooled to a rotational temperature of ∼3–10 K.